\begin{thebibliography}{34}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{{#1}}
\providecommand{\urlprefix}{URL }
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{DOI~\discretionary{}{}{}#1}\else
  \providecommand{\doi}{DOI~\discretionary{}{}{}\begingroup
  \urlstyle{rm}\Url}\fi
\providecommand{\eprint}[2][]{\url{#2}}

\bibitem[{Anderson(2001)}]{Anderson2001}
Anderson R (2001) {Why Information Security is Hard - An Economic Perspective}.
  In: ACSAC

\bibitem[{Anderson et~al(2012)Anderson, Barton, Rainer, Clayton, van Eeten,
  Levi, Moore, and Savage}]{Anderson2012}
Anderson R, Barton C, Rainer B, Clayton R, van Eeten M, Levi M, Moore T, Savage
  S (2012) {Measuring the Cost of Cybercrime}. In: WEIS

\bibitem[{Asghari et~al(2015{\natexlab{a}})Asghari, Ciere, and {Van
  Eeten}}]{Asghari2015a}
Asghari H, Ciere M, {Van Eeten} MJG (2015{\natexlab{a}}) {Post-Mortem of a
  Zombie: Conficker Cleanup After Six Years}. In: USENIX Security

\bibitem[{Asghari et~al(2015{\natexlab{b}})Asghari, van Eeten, and
  Bauer}]{Asghari2015b}
Asghari H, van Eeten MJ, Bauer JM (2015{\natexlab{b}}) {Economics of Fighting
  Botnets: Lessons from a Decade of Mitigation}. IEEE S{\&}P 13(5)

\bibitem[{Bilge et~al(2011)Bilge, Kirda, Kruegel, and Balduzzi}]{Bilge2011}
Bilge L, Kirda E, Kruegel C, Balduzzi M (2011) {EXPOSURE: Finding Malicious
  Domains Using Passive DNS Analysis}. In: NDSS

\bibitem[{Canali et~al(2011)Canali, Cova, Vigna, and Kruegel}]{Canali2011}
Canali D, Cova M, Vigna G, Kruegel C (2011) {Prophiler: A Fast Filter for the
  Large-Scale Detection of Malicious Web Pages}. In: WWW

\bibitem[{{\c{C}}etin et~al(2018){\c{C}}etin, Ga{\~{n}}{\'{a}}n, Altena,
  Tajalizadehkhoob, and van Eeten}]{CetinGATE18}
{\c{C}}etin O, Ga{\~{n}}{\'{a}}n C, Altena L, Tajalizadehkhoob S, van Eeten M
  (2018) Let me out! evaluating the effectiveness of quarantining compromised
  users in walled gardens. In: USENIX SOUPS

\bibitem[{{\c{C}}etin et~al(2019){\c{C}}etin, Ga{\~{n}}{\'{a}}n, Altena,
  Kasama, Inoue, Tamiya, Tie, Yoshioka, and van Eeten}]{CetinGAKITTYE19}
{\c{C}}etin O, Ga{\~{n}}{\'{a}}n C, Altena L, Kasama T, Inoue D, Tamiya K, Tie
  Y, Yoshioka K, van Eeten M (2019) {Cleaning Up the Internet of Evil Things:
  Real-World Evidence on {ISP} and Consumer Efforts to Remove Mirai}. In: NDSS

\bibitem[{van Eeten et~al(2010)van Eeten, Bauer, Asghari, Tabatabaie, and
  Rand}]{Eeten2010}
van Eeten M, Bauer JM, Asghari H, Tabatabaie S, Rand D (2010) {The Role of
  Internet Service Providers in Botnet Mitigation: an Empirical Analysis based
  on Spam Data}. In: WEIS

\bibitem[{van Eeten et~al(2016)van Eeten, Lone, Moura, Asghari, and
  Korczy\'nski}]{EetenLMAK16}
van Eeten M, Lone Q, Moura G, Asghari H, Korczy\'nski M (2016) {Evaluating the
  Impact of AbuseHUB on Botnet Mitigation}. CoRR

\bibitem[{Felegyhazi et~al(2010)Felegyhazi, Kreibich, and
  Paxson}]{Felegyhazi2010a}
Felegyhazi M, Kreibich C, Paxson V (2010) {On the Potential of Proactive Domain
  Blacklisting}. In: USENIX LEET

\bibitem[{Fryer et~al(2015)Fryer, Stalla-Bourdillon, and Chown}]{Fryer2015}
Fryer H, Stalla-Bourdillon S, Chown T (2015) {Malicious web pages: What if
  hosting providers could actually do something}. Computer Law {\&} Security
  Review 31(4)

\bibitem[{Hao et~al(2011)Hao, Tech, Feamster, and Tech}]{Hao2011}
Hao S, Tech G, Feamster N, Tech G (2011) {Monitoring the Initial DNS Behavior
  of Malicious Domains}. In: ACM IMC

\bibitem[{Hao et~al(2013)Hao, Thomas, Paxson, Feamster, Kreibich, Grier, and
  Hollenbeck}]{Hao2013}
Hao S, Thomas M, Paxson V, Feamster N, Kreibich C, Grier C, Hollenbeck S (2013)
  {Understanding the Domain Registration Behavior of Spammers}. In: ACM IMC

\bibitem[{Korczy\'nski et~al(2017{\natexlab{a}})Korczy\'nski, Tajalizadehkhoob,
  Noroozian, Wullink, Hesselman, and van Eeten}]{tld_maciej}
Korczy\'nski M, Tajalizadehkhoob S, Noroozian A, Wullink M, Hesselman C, van
  Eeten M (2017{\natexlab{a}}) {Reputation Metrics Design to Improve
  Intermediary Incentives for Security of TLDs}. In: IEEE Euro S\&P

\bibitem[{Korczy\'nski et~al(2017{\natexlab{b}})Korczy\'nski, Wullink,
  Tajalizadehkhoob, Moura, and Hesselman}]{sadag}
Korczy\'nski M, Wullink M, Tajalizadehkhoob S, Moura GC, Hesselman C
  (2017{\natexlab{b}}) {Statistical Analysis of DNS Abuse in gTLDs Final
  Report}. Tech. rep.

\bibitem[{Korczy{\'{n}}ski et~al(2018)Korczy{\'{n}}ski, Wullink,
  Tajalizadehkhoob, Moura, Noroozian, Bagley, and Hesselman}]{Korczynski2018}
Korczy{\'{n}}ski M, Wullink M, Tajalizadehkhoob S, Moura GC, Noroozian A,
  Bagley D, Hesselman C (2018) {Cybercrime After the Sunrise: A Statistical
  Analysis of DNS Abuse in New gTLDs}. In: ACM ASIACCS

\bibitem[{Korczy\'nski et~al(2020)Korczy\'nski, Nosyk, Lone, Skwarek, Jonglez,
  and Duda}]{closed}
Korczy\'nski M, Nosyk Y, Lone Q, Skwarek M, Jonglez B, Duda A (2020) {The
  Closed Resolver Project: Measuring the Deployment of Source Address
  Validation of Inbound Traffic}. CoRR

\bibitem[{Le~Pochat et~al(2020)Le~Pochat, Van~hamme, Maroofi, van Goethem,
  Preuveneers, Duda, Joosen, and Korczy\'nski}]{PochathMGPDJK20}
Le~Pochat V, Van~hamme T, Maroofi S, van Goethem T, Preuveneers D, Duda A,
  Joosen W, Korczy\'nski M (2020) A practical approach for taking down
  avalanche botnets under real-world constraints. In: NDSS

\bibitem[{Levchenko et~al(2011)Levchenko, Pitsillidis, Chachra, Enright,
  Felegyhazi, Grier, Halvorson, Kanich, Kreibich, Liu, McCoy, Weaver, Paxson,
  Voelker, and Savage}]{Levchenko2011}
Levchenko K, Pitsillidis A, Chachra N, Enright B, Felegyhazi M, Grier C,
  Halvorson T, Kanich C, Kreibich C, Liu H, McCoy D, Weaver N, Paxson V,
  Voelker GM, Savage S (2011) {Click Trajectories: End-to-End Analysis of the
  Spam Value Chain}. In: IEEE S{\&}P

\bibitem[{Liu et~al(2011)Liu, Levchenko, F{\'{e}}legyh{\'{a}}zi, Kreibich,
  Maier, Voelker, and Savage}]{Liu2011}
Liu H, Levchenko K, F{\'{e}}legyh{\'{a}}zi M, Kreibich C, Maier G, Voelker GM,
  Savage S (2011) {On the Effects of Registrar Level Intervention}. In: USENIX
  LEET

\bibitem[{Lone et~al(2017)Lone, Luckie, Korczy\'{n}ski, and van Eeten}]{loops}
Lone Q, Luckie M, Korczy\'{n}ski M, van Eeten M (2017) {Using Loops Observed in
  Traceroute to Infer the Ability to Spoof}. In: PAM

\bibitem[{Luckie et~al(2019)Luckie, Beverly, Koga, Keys, Kroll, and
  kc~claffy}]{LuckieBKKKc19}
Luckie MJ, Beverly R, Koga R, Keys K, Kroll JA, kc~claffy (2019) {Network
  Hygiene, Incentives, and Regulation: Deployment of Source Address Validation
  in the Internet}. In: ACM CCS

\bibitem[{Maroofi et~al(2020)Maroofi, Korczy\'nski, Hesselman, Ampeau, and
  Duda}]{comar}
Maroofi S, Korczy\'nski M, Hesselman C, Ampeau B, Duda A (2020) {COMAR:
  Classification of Compromised versus Maliciously Registered Domains}. In:
  IEEE Euro S\&P

\bibitem[{Moura et~al(2015)Moura, Ga{\~{n}}{\'{a}}n, Lone, Poursaied, Asghari,
  and van Eeten}]{MouraGLPAE15}
Moura GCM, Ga{\~{n}}{\'{a}}n C, Lone Q, Poursaied P, Asghari H, van Eeten M
  (2015) {How Dynamic is the ISPs Address Space? Towards Internet-Wide {DHCP}
  Churn Estimation}. In: Networking

\bibitem[{Nikiforakis et~al(2011)Nikiforakis, Joosen, and
  Johns}]{Nikiforakis2011}
Nikiforakis N, Joosen W, Johns M (2011) {Abusing Locality in Shared Web
  Hosting}. In: EUROSEC

\bibitem[{Noroozian(2020)}]{Noroozian2020a}
Noroozian A (2020) {Evaluating Hosting Provider Security - Through Abuse Data
  and the Creation of Metrics}. Doctorol thesis, TU Delft, Faculty of
  Technology, Policy and Management

\bibitem[{Noroozian et~al(2015)Noroozian, Korczy\'nski, Tajalizadehkhoob, and
  van Eeten}]{Noroozian2015}
Noroozian A, Korczy\'nski M, Tajalizadehkhoob S, van Eeten M (2015) {Developing
  Security Reputation Metrics for Hosting Providers}. In: USENIX CSET

\bibitem[{Noroozian et~al(2017)Noroozian, Ciere, Korczy\'nski,
  Tajalizadehkhoob, and Eeten}]{Noroozian2017}
Noroozian A, Ciere M, Korczy\'nski M, Tajalizadehkhoob S, Eeten MV (2017)
  {Inferring the Security Performance of Providers from Noisy and Heterogenous
  Abuse Datasets}. In: WEIS

\bibitem[{Noroozian et~al(2019)Noroozian, Koenders, van Veldhuizen,
  Ga{\~{n}}{\'{a}}n, Alrwais, McCoy, and van Eeten}]{Noroozian2019}
Noroozian A, Koenders J, van Veldhuizen E, Ga{\~{n}}{\'{a}}n C, Alrwais S,
  McCoy D, van Eeten M (2019) {Platforms in Everything: Analyzing Ground-Truth
  Data on the Anatomy and Economics of Bullet Proof Hosting}. In: USENIX
  Security Symposium

\bibitem[{Stone-Gross et~al(2011)Stone-Gross, Holz, Stringhini, and
  Vigna}]{Stone-Gross2011a}
Stone-Gross B, Holz T, Stringhini G, Vigna G (2011) {The Underground Economy of
  Spam: A Botmaster's Perspective of Coordinating Large-Scale Spam Campaigns}.
  In: USENIX LEET

\bibitem[{Szurdi et~al(2014)Szurdi, Kocso, Cseh, Felegyhazi, and
  Kanich}]{Szurdi2014}
Szurdi J, Kocso B, Cseh G, Felegyhazi M, Kanich C (2014) {The Long Taile of
  Typosquatting Domain Names}. USENIX Security Symposium

\bibitem[{Tajalizadehkhoob et~al(2016)Tajalizadehkhoob, Korczy\'nski,
  Noroozian, Ga{\~{n}}{\'{a}}n, and van Eeten}]{Tajalizadehkhoobnoms}
Tajalizadehkhoob S, Korczy\'nski M, Noroozian A, Ga{\~{n}}{\'{a}}n C, van Eeten
  M (2016) {Apples, Oranges and Hosting Providers: Heterogeneity and Security
  in the Hosting Market}. In: IEEE/IFIP NOMS

\bibitem[{Tajalizadehkhoob et~al(2018)Tajalizadehkhoob, B{\"{o}}hme,
  Ga{\~{n}}{\'{a}}n, Korczy\'nski, and van Eeten}]{Tajalizadehkhoobtoit}
Tajalizadehkhoob S, B{\"{o}}hme R, Ga{\~{n}}{\'{a}}n C, Korczy\'nski M, van
  Eeten M (2018) {Rotten Apples or Bad Harvest? What We Are Measuring When We
  Are Measuring Abuse}. {ACM TOIT} 18(4)

\end{thebibliography}

\documentclass[graybox, natbib, nosecnum, twocolumn]{svmult}
\bibpunct{(}{)}{;}{a}{}{,} 

\usepackage{mathptmx}       
\usepackage{helvet}         
\usepackage{courier}        
\usepackage{type1cm}        

\usepackage{makeidx}         
\usepackage{graphicx}        
\usepackage{multicol}        
\usepackage[bottom]{footmisc}
\usepackage[normalem]{ulem}	
\usepackage{hyperref}  
\usepackage{soul}   

\usepackage{multirow}
\usepackage{booktabs}
\usepackage{enumitem}
\usepackage{comment}

\begin{document}

\title*{Security Reputation Metrics}
\author{Maciej Korczy\'nski \thanks{Univ. Grenoble Alpes, CNRS, Grenoble
INP, LIG, France} and Arman Noroozian \thanks{Delft University of Technology, Netherlands}} 
\authorrunning{Korczy\'nski and Noroozian}
\institute{Maciej Korczy\'nski \at Univ. Grenoble Alpes, CNRS, Grenoble INP, LIG, France, \email{maciej.korczynski@univ-grenoble-alpes.fr}
\and Arman Noroozian \at Delft University of Technology, Netherlands \email{a.noroozian@tudelft.nl}}
\maketitle

\section{Definitions}
Security reputation metrics (\textit{aka.} security metrics) quantify the security levels of organization (e.g., hosting or Internet access providers) relative to comparable entities. 
They enable benchmarking and are essential tools for decision and policy-making in security, and may be used to govern and steer responsible parties towards investing in security when economic or other decision-making factors may drive them to do otherwise.

\section{Background}
We increasingly interact with online digital content, which 
relies on services provided by so-called \textit{Internet intermediaries}, among them Internet service providers
(ISPs), hosting providers, domain name registrars, search engines, payment providers, certification authorities, cloud service providers, social network operators and e-commerce suppliers. Simultaneously, miscreants have been abusing the products and infrastructure of such service providers towards cybercrime by compromising their security and using them in unintended ways. 

A wealth of research into cybercrime points to how cybercriminals misuse hosting services \citep{Nikiforakis2011,Noroozian2019}, domain names \citep{Felegyhazi2010a,Liu2011,Hao2011,Hao2013,Szurdi2014,PochathMGPDJK20}, DNS services \citep{Bilge2011,Canali2011} and mail servers \citep{Stone-Gross2011a,Levchenko2011} to name a few examples.

A typical situation with intermediary services is that their consumers are at an inherently disadvantaged knowledge position in which they don't know much about how secure these services are. In contrast the service providers themselves know much more. 
Without additional security-related information, users typically base their decision to subscribe to a particular service on more readily available information, for instance pricing in case of hosting, or available bandwidth in case of broadband ISP. 
In other words, it is difficult for other businesses, consumers and regulators to reliably assess how secure intermediary services are. 
This so-called \textit{information asymmetry} about the security of intermediary services, combined with the fact that typically, parties other than the intermediaries themselves bear the cost of the cybercrime enabled through their services \citep{Anderson2012}, leads to an erosion of their incentives to adequately invest in security. 
In other words their economic incentives are misaligned with security goals.

Thus, systematically comparing the security performance of digital services, and intermediaries, may help reduce the security information asymmetries from a consumer perspective. 
Security reputation metrics are essential to this end and may help in reducing cybercrime, which is as much a technical issue as a problem of economic incentives \citep{Anderson2001}. 
Various stakeholders such as network service providers \citep{Asghari2015b}, domain registries \citep{tld_maciej}, law enforcement agencies \citep{Noroozian2015}, and even policy makers \citep{sadag} employ security metrics to answer questions like which are the worst service providers and what actions should be taken to steer market driven economies towards improved security outcomes. 

Existing metrics typically compare security based on either (\textbf{i}) how frequently abuse incidents occur (or vulnerabilities are discovered), i.e. are based on counting the number of incidents (vulnerabilities), or (\textbf{ii}) based on how timely incidents are remediated once they have occurred. 
The number of maliciously registered domain names, compromised end-user machines, and machines running outdated software per service provider are examples of the former case. 
The amount of time required to remediate, block and remove phishing or malware spreading web-pages are examples of the second metric type. 

Count based metrics are typically normalized by estimates of the size of each intermediary's potential \textit{attack surface} to control for more exposed intermediaries that have higher probability of experiencing incidents. This enables apples-to-apples comparisons between intermediaries of various exposure. The number of advertised IP addresses by a hosting provider, or the number of domains that it hosts, for instance may be used to estimate a hosting provider's potential attack surface \citep{Noroozian2015,Tajalizadehkhoobnoms, Tajalizadehkhoobtoit}.

\section{Application}

The following subsections present examples of security reputation metrics for different types of providers and show how they are used by Internet stakeholders in reducing cybercrime and aligning economic incentives with better security.

\subsection{Hosting Providers}

Hosting providers are companies that provide servers via which customers can make content or services available on the Internet e.g. websites, email or even sharing of files. As with all services on the Internet, they are also abused for criminal purposes. 
Think of phishing sites, command-and-control servers for botnets, child sexual abuse material (CSAM), malware distribution, and spam servers.

In theory, hosting providers can mitigate or prevent the abuse of their infrastructure by following security best practices set forth by organizations like the Messaging, Malware and Mobile Anti-Abuse Working Group (M3AAWG). For example, by vetting customers, monitoring their infrastructure for signs of compromise or even running anti-virus software to name a few recommended practices. Yet following such advice remains voluntary, and there is considerable variation in how hosting providers choose to act when it comes to abuse and securing their infrastructure. 
Naturally then, some providers are abused more often than others.  

Therefore, the question of which hosting providers are better at securing their infrastructure, is one that may be answered through security metrics that quantify how effective each provider is at curbing abuse. 

A systematic approach to metrics development within the hosting market is presented in a study by \citeauthor{Noroozian2015}, which is scoped to the Netherlands and the result of a collaboration between several local authorities. The aim of the study being to answer the following question: Which are the worst hosting providers in Dutch jurisdiction? Metrics developed within this study were used to steer the Dutch hosting market towards more effective security practices via involving several local authorities \citep{Noroozian2015}.
Furthermore, \citeauthor{Noroozian2017}, have developed metrics for comparing the security of hosting providers globally \citep{Noroozian2017}, which are shown to have considerable predictive power. Other work discusses how multiple approaches to curbing abuse within the hosting market, from loose market-based approaches, to stricter regulation-based approaches, may benefit from the employment of security metrics \citep{Fryer2015, Noroozian2020a}.

\subsection{Top-Level Domains}

There exists little empirical information about the security of entire Top-Level Domains such as .com, .nl,  or .top.
\cite{tld_maciej} were first to present security metrics for this ecosystem and have measured their operational values.  
They compared entire TLDs against the rest of the market. However, they have explicitly distinguished
those metrics from the  
objective of measuring the security performance of the registry operators. The reason is that a TLD is not a single organization but
constitutes an entire `domain name ecosystem' of different
types of intermediaries such as domain registries, registrars, or hosting providers.

The follow-up study requested by ICANN, spanning the period up to the end of 2016, investigated the following research question: How do abuse rates in the new gTLDs (e.g., .top, .science, .bank, or .site) compare to legacy gTLDs (e.g., .com, .net, .edu, or .org), since the introduction of the new gTLD program in 2013?
To determine the distribution of abusive activities across the gTLDs, \citeauthor{Korczynski2018} have analyzed the number of reported domains from reputed URL and domain blacklists normalized by the size of their respective TLD, calculated as the number of 2\textsuperscript{nd}-level domain names present in a zone file for each gTLD. 

They have compared abuse rates separately for compromised and maliciously registered domains \citep{comar}.
Reputation metrics reflecting spam activity in the new and legacy gTLDs have revealed an interesting trend: miscreants seem to be switching from abusing legacy to new
gTLDs when it comes to maliciously registered spam domains. 
In the last quarter of 2016, new gTLDs collectively had approximately one order of magnitude higher rate of spam domains per 10,000
registrations compared to legacy gTLDs.
Moreover, as many as 15 most abused new gTLDs had more than 10\% of all registered domain names blacklisted by Spamhaus at the end of 2016.
Finally, as many as 51.5\%, 47.6\%, and 33.4\% of all .science, .stream, and .study new gTLDs, respectively, were maliciously registered by cybercriminals and blacklisted by Spamhaus.

ICANN has used the calculated reputation metrics to review the existing anti--abuse safeguards in new gTLDs, and to introduce more effective ones before an upcoming new gTLD rollout.

\subsection{Internet Service Providers}
A significant amount of scientific work addresses the role of Internet Service Providers (ISPs)---network access providers---in mitigating cybercrime, for instance, through botnet mitigation by employing security metrics \citep{Eeten2010,Asghari2015a,EetenLMAK16}. 
ISPs typically provide Internet connectivity to customers and thus are in a unique position to mitigate certain forms of cybercrime at their origin. 

ISPs should also follow security best-practices 
to mitigate abuse.
Examples of some security best-practices for ISPs include the use of walled-gardens to quarantine and isolate infected machines connected to the Internet \citep{CetinGATE18,CetinGAKITTYE19}, or deploying Source Address Validation, also known as BCP38, to prevent Distributed Denial-of-Service (DDoS) attacks from being launched via their infrastructure \citep{LuckieBKKKc19,closed,loops}.
Yet, again, the voluntary nature of implementing such best practices results in certain ISPs experiencing a higher level of abuse than others due to having laxer security practices. 

Differences among ISPs in mitigating botnet infections, for instance, have been quantified in several studies. 
\citeauthor{Eeten2010} , for example, found that just 50 ISPs account for over half of all spam sources suggesting concentrations of spambots within a few ISPs worldwide. 

A typical approach in such studies is to first process global or national datasets of botnet activity from available sinkholes 
and to extract IP addresses of infected end-user machines. 
The methods employed typically map each bot-infected IP address to an ISP and then counts the IP addresses seen in each ISP 
per day to account for IP churn \citep{MouraGLPAE15}.
Security metrics to compare botnet mitigation among ISPs are then calculated by dividing this count by the numbers of subscribers of each ISP, where larger number indicate less effective mitigation by the ISP. 
Such security metrics for botnet mitigation among ISPs have been employed successfully to incentivize ISPs, within the Netherlands for example, to subscribe to threat intelligence data feeds, and deal with bot infections within their networks \citep{EetenLMAK16}.

\section{Open Problems and Future Directions}
Empirical measurements and analysis of security indicators leading to reliable security reputation metrics has proven to be quite challenging. 
The challenge partly lies in limitations of data: e.g. coverage, measurement errors and biases that are invariably linked to metric limitations. 
For example, the construction of security metrics typically depends on 
chaotic internet operations data such WHOIS information that are error prone and incomplete, or on ever evolving dynamic BGP routing data for attributing security incidents to the responsible entities. 

The security incidents themselves are observed through various opaque abuse feeds with no clear documentation of their collection methodologies, accuracy, and biases. Abuse feeds contain various degrees of false positive incident information or carry biases that are not well documented or understood. 
Abuse feed coverage is also limited with an unknown number of security incidents goes unnoticed as false negatives for each entity thereby affecting metric outcomes. 

In addition to limitations in data, methodological challenges in constructing metrics also exist. 
A particularly challenging methodological aspect is that of identifying service providers within certain markets. For example, in the case of hosting providers, there is no maintained authoritative list of companies, even at a country level, to identify the companies that offer hosting services. This problem is worsened by layers of smaller companies that resell the services of larger hosting providers.
Similar problems exist for domain registrars. Yet, such information is vital for the construction and comparison of service providers against their competitors at a market level to make security metrics more useful. 

Another methodological challenge is an incomplete causal understanding of the factors that drive abuse across online services. A better and more complete causal understanding of such drivers enriches and allows for the construction of security metrics that are better interpretable, easier to understand, and more useful.

Finally, there are practical limitations as well including the fact that metrics do not reflect the intent of bad service providers whether it be negligent or criminal behavior for instance in the case of bullet-proof services that cater to cybercriminals with the promise of ignoring or delaying lawful take down requests.  
As such, security metrics can be gamed by coordinated criminals and thus there are limits to how they may be interpreted.

\bibliographystyle{spbasic}  
\bibliography{SRM.bib}

\end{document}